# Post Hoc Regression Refinement via Pairwise Rankings

### v1

@@ 1 Introduction
Accurate prediction of continuous properties is crucial across scientific and engineering disciplines. In molecular-property prediction (MPP), for instance, reliable estimates of physical or chemical attributes can accelerate drug discovery, materials design, and catalyst development. Recent advances in deep learning have enhanced regression models by leveraging large datasets to uncover complex, nonlinear relationships between structured inputs (e.g., molecular graphs, crystal structures) and their associated properties. Yet, unlike computer vision or natural-language processing, where large labeled corpora can be mined or crowd-sourced, many specialized fields face a fundamental bottleneck: acquiring ground-truth labels requires expert-led experiments that are both costly and slow. Consequently, real-world tasks often operate in data-scarce regimes where even 50 labeled samples may be considered plentiful [Sun et al., 2024].
@@
We introduce RankRefine, a post-hoc refinement framework that improves regression predictions by incorporating pairwise rankings. Given a query sample and a small set of labeled references, which can conveniently be drawn from training data, a ranker infers the relative ordering between the query and each reference; these comparisons define a likelihood whose minimization yields a rank-based property estimate. RankRefine then fuses this estimate with the base regressor’s prediction via inverse-variance weighting, all without retraining.
Because RankRefine is largely model-agnostic and requires no architectural changes, it suits low-data, few-shot, and meta-learning scenarios. Leveraging ranking signals from publicly hosted LLMs, it provides substantial improvements in regression accuracy with minimal labeled data and negligible client-side computation. Experiments on synthetic and real-world benchmarks, including multiple MPP tasks, demonstrate that RankRefine consistently boosts predictive performance. Further analysis links its effectiveness to ranking quality, underscoring the promise of LLMs and other expert-informed rankers as practical complements to data-scarce regression models. The source code is provided in the supplementary materials.
@@ 4.8 Effects of Noisy Variance Estimates
We study the effect of noise in the ranker variance estimate on the Solubility dataset using $k=30$ and $k=40$. We add uniform perturbations sampled from [-b, +b] to the estimated ranker variance before fusion. As shown in Figure 6, performance degradation becomes noticeable when $b\geq 5$, which is more than three times the standard deviations of the ranker variance estimates. For reference, the mean and standard deviation of the estimated ranker variance are 2.358 ± 1.079.
@@
Figure 6: Impact of noisy ranker variance estimate on the regression performance in Caco2 dataset. Uniform noise from [-b, +b] is added to the estimated ranker variance. Performance starts to degrade when $b\geq 5$, more than 3 times the standard deviation of the ranker variance estimate.
@@ B.1 Prompt for obtaining pairwise rankings from ChatGPT-4o
* Before you start deisgning the heuristics, briefly explain the property

### v2

@@ 1 Introduction
Accurate prediction of continuous properties is crucial across scientific and engineering disciplines. In molecular‐property prediction (MPP), for instance, reliable estimates of physical or chemical attributes can accelerate drug discovery, materials design, and catalyst development. Recent advances in deep learning have enhanced regression models by leveraging large datasets to uncover complex, nonlinear relationships between structured inputs (e.g., molecular graphs, crystal structures) and their associated properties. Yet, unlike computer vision or natural-language processing, where large labeled corpora can be mined or crowd-sourced, many specialized fields face a fundamental bottleneck: acquiring ground-truth labels requires expert-led experiments that are both costly and slow. Consequently, real-world tasks often operate in data-scarce regimes where even 50 labeled samples may be considered plentiful [Sun et al., 2024].
@@
We introduce RankRefine, a post hoc refinement framework that improves regression predictions by incorporating pairwise rankings. Given a query sample and a small set of labeled references, which can conveniently be drawn from training data, a ranker infers the relative ordering between the query and each reference; these comparisons define a likelihood whose minimization yields a rank-based property estimate. RankRefine then fuses this estimate with the base regressor’s prediction via inverse-variance weighting, all without retraining.
Because RankRefine is largely model-agnostic and requires no architectural changes, it suits low-data, few-shot, and meta-learning scenarios. Leveraging ranking signals from publicly hosted LLMs, it provides substantial improvements in regression accuracy with minimal labeled data and negligible client-side computation. Experiments on synthetic and real-world benchmarks, including multiple MPP tasks, demonstrate that RankRefine consistently boosts predictive performance. Further analysis links its effectiveness to ranking quality, underscoring the promise of LLMs and other expert-informed rankers as practical complements to data-scarce regression models. The source code is available at https://github.com/ktirta/regref.
@@ 4.6 Few-Shot Molecular-Property Prediction with LLMs
One natural concern when employing a pretrained LLM as a ranker is whether it simply memorizes from training data which item (e.g., a molecule) has a higher or lower property value (e.g., solubility) compared to others. To test this, we conducted an additional evaluation on a private compound-activity dataset consisting of 75 compounds with IC50 labels measured against a specific target protein complex. This dataset is unpublished, and therefore, not part of any LLM training corpora, ruling out direct memorization. On this dataset, ChatGPT-4o achieves 60.14% pairwise ranking accuracy, indicating generalization capability and not simple memorization.
@@ 4.8 Effects of Noisy Variance Estimates
We study the effect of noise in the ranker variance estimate on the Solubility dataset using $k=30$ and $k=40$. We add uniform perturbations sampled from [-b, +b] to the estimated ranker variance before fusion. As shown in Figure 6, performance degradation becomes noticeable when $b\geq 5$, which is more than three times the standard deviation of the ranker variance estimates. For reference, the mean and standard deviation of the estimated ranker variance are 2.358 ± 1.079.
@@
Figure 6: Impact of noisy ranker variance estimate on the regression performance in Solubility dataset. Uniform noise from [-b, +b] is added to the estimated ranker variance. Performance starts to degrade when $b\geq 5$, more than 3 times the standard deviation of the ranker variance estimate.
4.9 Effects of Biased Regressor
We systematically inject constant bias into an otherwise unbiased regressor. Specifically, we uniformly draw ground-truth labels $y_{0}\sim U(0,1)$, and set predictions by injecting zero-mean Gaussian noise $\hat{y}_{0}^{reg}=y_{0}+\epsilon$, such that $\text{MAE}\approx 1\text{SD}$ where SD is the standard deviation of the ground-truth labels. We then add a constant offset, $B$, increasing in magnitude from 0% to 70% of the label standard deviation (SD). We report the refined-to-original MAE ratio, $\beta$, for varying ranker accuracies. Table 3 in Appendix D shows that $\beta$ worsens as bias increases, but remains <<1 up to $B$ = 60% SD, indicating that RankRefine remains effective even under substantial regressor bias.
4.10 Effects of Biased Sampling
We study the effects of non-uniform sampling of references, in particular, when the references are clustered. Similar to the biased regressor experiment, we generate synthetic predictions with ground-truth label range $[0,1]$. We then restrict the reference labels to increasingly narrow subranges of the ground-truth label range, centered around the midpoint of the range. The reference set bias, $RB=x\%$, means that the reference labels cover only $(100-x)\%$ of the full ground-truth range. For example, an $RB$ = 10% means the reference labels range is [0.05, 0.95]. Table 4 in Appendix E suggests that RankRefine performance is best when the reference set spans the full ground-truth labels’ range (i.e., $RB$ = 0%), but clustered sets still yield substantial gains. For example, at RB = 90% and ranker accuracy = 60%, $\beta$ = 0.8842, an 11.58% improvement.
Furthermore, we evaluate biased sampling in two challenging settings: (i) disjoint: reference labels are in [-1,0] or [1,2] while queries are in [0,1]; and (ii) partial overlap: reference labels are in [-0.5, 0.5] or [0.5, 1.5], while queries are in [0,1]. This simulates a distribution shift, where the distribution of the reference labels is different from the distribution of the query labels. Despite the challenging setups, RankRefine improves the base regressor when the ranker is moderately accurate (i.e., $\geq$65% in the disjoint case, $\geq$55% in the partial overlap case). More details are available in Table 5 in Appendix E.
@@ B.1 Prompt for obtaining pairwise rankings from ChatGPT-4o
* Before you start designing the heuristics, briefly explain the property
@@ Appendix C More Comparisons with the Projection-based Approach
Appendix D Ablation: Biased Regressor
We inject constant bias $B$ into an otherwise unbiased regressor, where the magnitude of the offset is measured as a percentage of the ground truth labels standard deviation. RankRefine offers improvement even when the regressor is biased (e.g., 60% of the SD of the ground truth labels)
Table 3: RankRefine performance, measured in $\beta$ (lower is better), on a biased regressor.
Ranker Accuracy 50% 55% 60% 65% 70%
Bias (B) = 0% SD 0.8911 0.8436 0.8694 0.8668 0.8949
B = 10% SD 0.8995 0.8800 0.8623 0.8602 0.8969
B = 20% SD 0.9084 0.8869 0.8704 0.8704 0.9047
B = 30% SD 0.9224 0.8997 0.8844 0.8861 0.9192
B = 40% SD 0.9409 0.9172 0.9059 0.9071 0.9395
B = 50% SD 0.9645 0.9403 0.9329 0.9343 0.9654
B = 60% SD 0.9942 0.9702 0.9649 0.9670 0.9967
B = 70% SD 1.0291 1.0062 1.0022 1.0049 1.0326
Appendix E Ablation: Biased Sampling
We restrict the reference set range to increasingly narrow subsets of the ground truth label range ([0, 1]), simulating clustered or biased samples. The reference set bias, $RB=x\%$, means that the reference labels cover only $(100-x)\%$ of the full ground-truth range. For example, an $RB$ = 10% means the reference labels range is [0.05, 0.95]. Best performance is achieved when the reference set spans the entire range of the ground truth labels, but clustered reference sets still provide substantial regression improvements.
Table 4: RankRefine performance, measured in $\beta$, with non-uniform sampling of the reference set.
Ranker Accuracy 50% 55% 60% 65% 70%
Reference Set Bias (RB) = 0% range 0.8828 0.8706 0.8343 0.8443 0.8598
RB = 10% range 0.9034 0.8731 0.8493 0.8529 0.8768
RB = 20% range 0.9078 0.8753 0.8366 0.8694 0.9002
RB = 30% range 0.9076 0.8701 0.8439 0.8733 0.9280
RB = 40% range 0.9143 0.8767 0.8618 0.8824 0.9414
RB = 50% range 0.9253 0.8786 0.8640 0.8769 0.9480
RB = 60% range 0.9349 0.8858 0.8712 0.8929 0.9546
RB = 70% range 0.9385 0.8919 0.8770 0.8957 0.9597
RB = 80% range 0.9551 0.8984 0.8798 0.8943 0.9558
RB = 90% range 0.9664 0.9042 0.8842 0.9035 0.9631
To test the impact of distribution shift between reference and query, we created two challenging scenarios where the reference labels and query labels are drawn uniformly from different range values:
  • Disjoint distributions: reference labels in [-1, 0] or [1, 2], query labels in [0, 1]. We averaged the results of the two cases, that is, between ref = [-1, 0], query = [0, 1] and ref = [1, 2], query = [0, 1].
  • Partial overlap: reference labels in [-0.5, 0.5] or [0.5, 1.5], query labels in [0, 1]. We averaged the results of the two cases, that is, between ref = [-0.5, 0.5], query = [0, 1] and ref = [0.5, 1.5], query = [0, 1].
The results in 5 shows that RankRefine still improves the base regressor when the ranker is moderately accurate, even when the distribution of the reference labels is different from the query.
Table 5: RankRefine performance, measured in $\beta$, for disjoint and partially overlapping reference set.
Ranker Accuracy 50% 55% 60% 65% 70%
Disjoint 1.2906 1.1629 1.0170 0.9437 0.9219
Partial Overlap 1.0160 0.9465 0.8700 0.8607 0.8898